shift summary: owl 9/23/99

John Shields (jshields@ksera)
Thu, 23 Sep 1999 07:04:07 -0500

Shift Summary for owl shift, 09/23/99

On Shift: J. Shields, S. Glazov

Hours of beam available at all intensities: 4.0
Hours of beam available at nominal intensity: 4.0
Hours of E799 data (special runs not included): 3.0
Hours of special runs (laser scans, muon runs...): 0.3

Accounting of lost time:

DAQ errors (kumquats, VV' triggers), bad tape put into drive
========================================================================
Runs ended this shift:
Run 14677 - pedestal run
Run 14680 - 6.3M events - BEAM_799

Tapes brought to FCC this shift:
(including tapes for online split)

UP*019-021, UPS007
========================================================================
Accesses during shift:

none

========================================================================
Overview of shift:
tevatron problems, VV' problems, DAQ problems, E799 datataking

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14677 00:35 begin pedestal run

14677 00:55 end ped run

01:10 MCR calls and promises beam. It comes in 5 min.

01:17 Crew unsuccessfully attempts to begin BEAM_799
run. Kumquat errors--> reset --> crate 53 full,
yellow light on spill counters. Stop and restart
new run.

01:35 all errors cleared, and KteV is ready to take
data! Unfortunately, beam dies.

14680 02:42 MCR calls & promises to deliver nice beam

02:42, 2 sec beam arrives

03:10 Looking at histograms, crew sees things that don't
look good-- VV' has a hot chanel. Valeri
courageously drags himself out of bed and visits
KteV to investigate. Verdict is that problem
needs an access to be fixed, so Valeri decides
to simply wait a few hours for the shutdown.

05:40 tapes full, run stopped. Tapes UP*019-021, UPS007
brought to FC

05:50 crew attempts to begin a new run, but (surprise!)
gets yet *another* "Error starting run" message.
Did DRT_FRESH_START & tried again. No dice. New
Error msg is "ftt error at KTEV3, drive 2. Crew
cleans drive. Doesn't help. Switched tape
to new drive & error moves to new drive. Net
result: tape was bad. Shift Crew is *not* amused.

14685 06:25 now that the tape mystery is solved KteV is once
again rarin' ta take data!

06:26 ...And once again, RIGHT on cue, beam vanishes!
Ch 13 message is "TeV PS in bypass, investigating...."
Again, shift crew is NOT amused.

06:49 MCR calls and confirms crew's deep-seeded
suspicion that "investigation" would be longer
than the 34 minutes until shutdown. Crew morale
drifts from non-amusement, to depression, and then
finally comes to rest in tired apathy.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift